the perimeter of the triangle is 218 inches. Find the length of each. 2x inches, x inches, (6x-7) inches
Answer by dragonwalker(73)

I assume each side is above so basically add each side together and that will equal 218 in then solve for x:
2x + x + 6x-7 = 218
rearrange so all X's on one side and numbers on the other. If you move one from one side of the = to the other side then change the sign so:
2x + x + 6x = 218 + 7
9x = 225
so divide both sides by 9 to find x:
9x/9 = 225/9
x = 25
so x = 25
Now we can find the value of each side.
First one was 2x so 2 times 25 = 50
The second one was x so is 25
The third was 6x-7 = 6(25) - 7
= 143
To check add the three sides together to see if they equal 218 so:
50 + 25 + 143 = 218!!!